Johnny Appleseed inherits $25,000 from his uncle and decides to invest part of it in a mutual fund which pays 7.2% annual interest and the remainder in a savings bond that pays 4.5% annual interest. If Johnny's annual interest totals $1525.95, how much did he invest in each account?
Answer by checkley77(12569)

.072x+.045(25,000-x)=1,525.95
.072x+1,125-.045x=1,525.95
.027x=1,525.95-1,125
.027x=400.95
x=400.95/.027
x=14,850 answer for amount invested @ 7.2%
25,000-14,850=10,150 invested @4.5%
Proof:
.072*14,850+.045*10,150=1,525.95
1,069.2+456.75=1,525.95
1,525.95=1,525.95